OpenStudy (lexipoo1998):

Can someone help me with a word problem? The number of milligrams of a drug that remains in a patient’s system after t hours is given by the function A(t) = Iert. Don was given 300 milligrams of medicine which leaves his bloodstream at a rate of 20%. How much of the medicine remains in his system after 4 hours?

OpenStudy (lexipoo1998):

@Luigi0210 @agent0smith Can you guys help?

OpenStudy (cggurumanjunath):

A(t) = Iert what is LERT ?

OpenStudy (lexipoo1998):

*IE^rt

OpenStudy (lexipoo1998):

@Euler271 @ChristopherToni

OpenStudy (lexipoo1998):

Can anyone please help me out?

OpenStudy (lexipoo1998):

The answer choices are a. 95.87 mg b.12.52 mg c.134.80 mg d.31.63 mg

OpenStudy (bibby):

Start by identifying the values you have. l (not sure) r (rate) t (time)

OpenStudy (lexipoo1998):

Umm I am not sure but I think it is Ie(capital i not l) u am really lost

OpenStudy (bibby):

e is a constant (2.2718..) and I assume l is the dosage

OpenStudy (lexipoo1998):

Oh... wait so how would I solve it? Can you or @Mertsj walk me through it?

OpenStudy (mertsj):

I think you have posted the formula incorrectly.

OpenStudy (lexipoo1998):

It says A(t) = Ie^rt so now I really do not know..

OpenStudy (mertsj):

I'm pretty sure the formula is a(t)=Ie^(-rt) but have it your way.

OpenStudy (bibby):

Don was given 300 milligrams of medicine which leaves his bloodstream at a rate of 20%. 4 hours. You clearly have the rate, doseage, e, and the time

OpenStudy (lexipoo1998):

If you can solve it with that then please do @Mertsj . This is the last question in my workbook packet and i just want to be able to solve it. Can you solve it?

OpenStudy (lexipoo1998):

So do i just mutlply it all out @bibby

OpenStudy (mertsj):

I cannot solve it using the formula you posted.

OpenStudy (lexipoo1998):

Can you solve it using something else which will result in one of the answer choices i have been given?

OpenStudy (lexipoo1998):

Well when i multiply it all out i didn't get an answer choice sooo @bibby what do i do?

OpenStudy (mertsj):

Yes. I can solve it using the correct formula which you refuse to post.

OpenStudy (lexipoo1998):

Alright so this is the formula: a(t)=Ie^(-rt) Can you please solve it now @Mertsj

OpenStudy (bibby):

I solved it using mertsj's negative version. You're not multiplying all the way. a(t)=Ie^(-rt)

OpenStudy (lexipoo1998):

Did you get 134.80?

OpenStudy (bibby):

Yep

OpenStudy (lexipoo1998):

Ok thank you so much!